Circular No. 9125

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 D4 (WISE) A. Mainzer, Jet Propulsion Laboratory, notes that an object found

on WISE satellite images (discovery observation tabulated below) appears “soft” in 12-μm images with a ‘blurry’ central condensation of size ∼ 14′′; 22-μm images show it to be brighter, with diameter ∼ 17′′. Numerous ground-based observers have been unable to detect cometary activity.

2010 UT α_{2000} δ_{2000} Observer
Feb. 28.01782 17^{h}15^{m}03.^{s}30 +41^{o}02′22′′.1 WISE

The available astrometry, parabolic orbital elements (T = 2009 Mar. 30.939 TT, q = 7.17413 AU, ω = 43^{o}.916, Ω = 266^{o}.488, i = 105^{o}.482, equinox 2000.0), and an ephemeris appear on MPEC 2010-E63.

COMET P/2010 E2 (JARNAC) An apparently asteroidal object reported by T. Glinos on CCD images

taken with a 0.64-m f/7.2 Ritchey-Chretien telescope at the Jarnac Observatory in Vail, AZ, U.S.A., on Mar. 9, 10, and 12 (discovery observation tabulated below; observers listed as D. Levy, W. Levy, and T. Glinos), and linked together by the Minor Planet Center staff and also with observations from the Mount Lemmon survey taken on Feb. 17 before being posted on the MPC’s ‘NEOCP’ webpage, has been found to show cometary appearance by other CCD astrometrists. D. Chestnov and A. Novichonok write that three co-added 300-s unfiltered CCD images taken remotely with a 0.36-m f/3.8 reflector at the Tzec Maun Observatory near Mayhill, NM, U.S.A. on Mar. 13.23 UT show a faint 0′.3 coma of mag 19 with a strong nuclear condensation and no tail. CCD images taken by W. H. Ryan and E. V. Ryan with the Magdalena Ridge Observatory (not McDonald Observatory, as erroneously written on IAUC 9117) 2.4-m f/8.9 reflector on Mar. 13.4 show a coma extending toward p.a. ∼ 280^{o}.

2010 UT  α_{2000}  δ_{2000}  Mag.
Mar. 9.30939  12^{h}02^{m}57.^{s}94  −1^{o}17′03′′.0  18.7

The available astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2010-E64.

T = 2010 Apr. 7.8086 TT  ω = 8.^{o}2567
e = 0.721879  Ω = 177.8984 }2000.0
q = 2.397960 AU  i = 15.4256
a = 8.621986 AU n^{o} = 0.0389308 P = 25.3 years
2010 March 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green